OpenStudy (anonymous):

cos(theta+phi)cos(theta-phi)=cos^2 theta-sin^2 phi can anyone can prove this

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

@iGreen ,@rosedewittbukater ,@frizzank Help me please

OpenStudy (anonymous):

@rosedewittbukater

OpenStudy (anonymous):

@frizzank

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

\[LHS\] \[= \cos(\theta + \phi). \cos(\theta - \phi)\] \[= \frac{1}{2} (2.\cos(\theta + \phi). \cos(\theta - \phi))\] \[= \frac{1}{2} (\cos(\theta + \phi + \theta - \phi). \cos(\theta + \phi - \theta + \phi))\] \[= \frac{1}{2} (\cos~2 \theta + \cos~2 \phi)\] \[= \frac{1}{2} (2 \cos^2 \theta - 1 + (1 - 2 \sin^2 \phi))\] \[= \frac{1}{2} (2 \cos^2 \theta - 1 + 1 - 2\sin^2 \phi)\] \[= \frac{1}{2} ~2(\cos^2 \theta - \sin^2 \phi)\] \[= \cos^2 \theta - \sin^2 \phi\] \[= RHS\] \[Q.E.D\]